OpenStudy (anonymous):

write the first four terms of the sequence for n=1,2,3, and 4. f(n)=2n^2

OpenStudy (anonymous):

f(1) = 2*1² = 2 f(2) = 2*2² = 2*4 = 8 f(3) = 2*3² = 2*9 = 18 f(4) = 2*4² = 2*16 = 32

OpenStudy (lalaly):

2, 8, 18, 32

OpenStudy (anonymous):

thank u sooo... much